A vector has a magnitude of 8.1

8.1 and is in quadrant 1 with an angle of 40.2


40.20 with respect to the +y

+y-axis. What is the x

x-component of the vector?


1
Expert's answer
2022-03-11T08:27:57-0500
"a_y=a\cos\theta=8.1\cos40.2^\circ=6.2\\
a_x=a\sin\theta=8.1\sin40.2^\circ=5.2"